If the point (-3,1/2) lies on the graph of the equation 2x+ky=-11, find the value of k.
a. -5/2 b. -34 c. -17/2 d. -10 e. none of the above
i came across this and i dont remember??

Answer by ewatrrr(24785)


Hi,
2x + ky = -11 Using ordered pair Pt(-3,1/2) to solve for k
2*-3 + k(1/2) = -11
-6 + k/2 = -11 Multiplying each term by 2 so as all denominators = 1
-12 + k = -22
k = -10
y = (2/10)x + 11/10